A vending machine is designed to dispense a mean of 7.6 oz of coffee into an 8 oz cup. The standard deviation of the amount of coffee dispensed is .4 oz.
a. Find the percent the machine will dispense from 7.4 oz to 7.7 oz.
b. Dispense less than 6.9 oz
I have it figured out that from 7.6 to 7.2 on the bell shaped curve is 34% but when I tried to add it up I couldn't go any where. i also got a probable answer of 12.75% for a. Please help!

Answer by stanbon(75887)

A vending machine is designed to dispense a mean of 7.6 oz of coffee into an 8 oz cup. The standard deviation of the amount of coffee dispensed is .4 oz.
a. Find the percent the machine will dispense from 7.4 oz to 7.7 oz.
z(7.4) = (7.4-7.6)/0.4 = -0.5
z(7.7) = (7.7-7.6)/0.4 = 0.25
---
P(7.4< x < 7.7) = P(-0.5< z < 0.25) = normalcdf(-0.5,0.25) = 0.2902
------------------------------
b. Dispense less than 6.9 oz
z(6.9) = (6.9-7.6)/0.4 = -1.75
---
P(x < 6.9) = P(z < -1.75) = normalcdf(-100,-1.75) = 0.0401
----------------------
Cheers,
Stan H.